OpenStudy (anonymous):

Someone please help! Will give medal The attention span of a two year-old is normally distributed with a mean of about 9.5 minutes and a standard deviation of 1.05 minutes. Suppose we randomly survey 60 two year-olds. What is the probability that a two year old has an attention span between 9.25 and 9.5 minutes?

OpenStudy (anonymous):

9.5+1.05=what?

OpenStudy (anonymous):

go on,

OpenStudy (anonymous):

ok 9.5+1.05=10.55

OpenStudy (anonymous):

and,60+10.55=70.55

OpenStudy (anonymous):

9.25+9.5=18.75

OpenStudy (anonymous):

and so, 70.55+18.75=89.3 is ur answer

OpenStudy (anonymous):

medal plz?

OpenStudy (anonymous):

plz?

OpenStudy (anonymous):

Sorry i had stepped away from the computer! Thank you very much!